Exercise F Never Late Delivery currently delivers packages for $9 each. The variable cost is $3 per package, and fixed costs are $60,000 per month. Compute the break-even point in both sales dollars and units under each of the following independent assumptions. 1. The costs and selling price are as given above. 2. Fixed costs are increased to $75,000. 3. Selling price is increased by 10%. (Fixed costs are $60,000.) 4. Variable cost is increased to $4.50 per unit. (Fixed costs are S60,000 and selling price is $9.) 5. Calculate the margin of safety for each option assuming sales were $140,000. 6. Comment on why the break-even points are different and which option would be the best. Problem G a. Assume that fixed costs of Celtics Company are $180,000 per year, variable cost is $12 per unit, and selling price is $30 per unit. Determine the break-even point in sales units b. The sales of Niners Corporation last year amounted to $20,000,000, its variable costs were $6,000,000, and its fixed costs were $4,000,000. At what level of sales dollars would the Niners Corporation break even? c. Answer each of the following questions labeled 1-4 based on the Niners Corporation from part b. What would have been the net income of the Niners Corporation, if sales volume had been 10% higher but selling prices had remained unchanged? Use the contribution margin income statement format. What would have been the net income of the Niners Corporation, if variable costs had been 10% lower? Use the contribution margin income statement format. What would have been the net income of the Niners Corporation, if fixed costs had been 10% lower? Use the contribution margin income statement format. Determine the break-even point in sales dollars for the Niners Corporation on the basis of the data given in (2) and then in (3). 1. 2. 3. 4.
Answer #1
Exercise F
1 Selling Price per unit $9
Variable costs per unit $3
Contribution margin per unit $6 0.666667
Fixed Costs $60,000
Break Even point ($ ) = 60000/0.67 90000
Break Even Point (Units) = 60000/6 10000
2 Fixed Costs $75,000
Break Even point ($ ) = 75000/0.67 112500
Break Even Point (Units) = 75000/6 12500
3 Selling price 9*110% $9.9
Variable costs per unit $3.0
Contribution margin per unit $6.9 0.696969697
Break Even point ($ ) = 60000/0.70 86086.95652
Break Even Point (Units) = 60000/6.9 8695.652174
4 Selling Price per unit $9.00
Variable costs per unit $4.50
Contribution margin per unit $4.50 0.5
Fixed Costs $60,000
Break Even point ($ ) = 60000/0.5 120000
Break Even Point (Units) = 60000/4.5 13333.33333
5 Option 1 Option 2 Option 3 Option 4
Sales $140,000 $140,000 $140,000 $140,000
Less : Break Even point Sales 90000 112500 86086.95652 120000
Margin of Safety $50,000 $27,500 $53,913 $20,000
6 The break even points are different due to change in contribution margin and fixed costs which are two component for computing break even point
Option 3 is preferred since it has lower break even point and higher margin of safety
Problem G
a Breakeven Point (Units ) = (180000/(30-12)) 10000
b Sales $20,000,000
Less : Variable Costs $6,000,000
Contribution Margin $14,000,000 0.7
Break Even Sales ($) = $4000000/(0.70) 5714285.714
Income Statement Option 1 Option 2 Option 3
Sales 22000000 $20,000,000 $20,000,000
Less : Variable Costs $6,000,000 $5,400,000 $6,000,000
Contribution Margin $16,000,000 $14,600,000 $14,000,000
Less : Fixed Costs $4,000,000 $4,000,000 $3,600,000
Net Income $12,000,000 $10,600,000 $10,400,000
Contribution Margin (%) 0.73 0.70
Break Even Sales ($) $5,479,452.05 $5,142,857.14
4000000/0.73 3600000/0.7